OpenStudy (anonymous):

Please help me?????? IN NEED OF HELP ASAP. Which of the following is the rational exponent expression of 4√ f?

OpenStudy (anonymous):

My answer is f^1/4

OpenStudy (anonymous):

Thats correct

OpenStudy (anonymous):

Is your root a 4th root, or is that like "4 multiplied by square root of f"?

OpenStudy (anonymous):

Thank you :) @ivanmlerner

OpenStudy (anonymous):

F to the 1/4th power

OpenStudy (anonymous):

\[\sqrt[4]{f} = f ^{\frac{ 1 }{ 4 }}\]

OpenStudy (anonymous):

But \[4\sqrt{f} = 4f ^{\frac{ 1 }{ 2 }}\]

OpenStudy (anonymous):

It depends on how that 4 is written... it's hard to tell in text equations like these.

OpenStudy (anonymous):



OpenStudy (anonymous):

no, I get that... that is correct for the problem of converting "4th root of f" to rational exponent form. From what you wrote, I just wasn't sure that was the correct problem here... "4√ f" looks to me like "4 times square root of f" It's up to you... you can see the actual problem. Just trying to help :)

OpenStudy (anonymous):

I'm sorry, I'm not mad in any form

OpenStudy (anonymous):

Was I correct nonetheless?

OpenStudy (anonymous):

No, I believe you :)

OpenStudy (anonymous):

I see what I did wrong now....! ooops ;)

OpenStudy (anonymous):

Do you get what I'm asking? I was trying to clarify your problem, not correct your answer.

OpenStudy (anonymous):

Yes, I'm very aware now, thank you for the heads up!

OpenStudy (anonymous):

You were correct... but for the wrong problem ;) Glad you caught it :)

OpenStudy (anonymous):

Oooopsie :)